OpenStudy (anonymous):

(8 x 10^5) / (4 x 10^2) in scientific notation.

Mehek (mehek14):

divide 8/4 = ?

OpenStudy (anonymous):

I'll show you my work

OpenStudy (anonymous):

(8 x 10^5) / (4 x 10^2) = (8 x 4) / (10^5 x 10 ^2) = 32 / 10^7

OpenStudy (anonymous):

then 3.2 / 10 ^6

Mehek (mehek14):

why would u multiply? / = divide so 8/4 = 2 now you would subtract the exponents \(\tt{\dfrac{10^5}{10^2}=10^{5-2}}\)

OpenStudy (anonymous):

oops i added

OpenStudy (anonymous):

so its 10 3

Mehek (mehek14):

\(\tt{2*10^3}\)

OpenStudy (anonymous):

thanks for pointing out my mistake really appreciated :)

Mehek (mehek14):

yw :)